You have a portfolio that is 37 percent invested in Stock R, 21 percent invested in Stock S, with the remainder in Stock T. The expected return on these stocks is 9.0 percent, 10.4 percent, and 12.7 percent, respectively. What is the expected return on the portfolio?
Answer #1

portfolio return=Respective return*Respective weight

=(0.37*9)+(0.21*10.4)+(0.42*12.7)(Remainder investment =(100-(37+21))=42%)

=10.848%